# MMVP: A Multimodal MoCap Dataset with Vision and Pressure Sensors

### v1

@@ Abstract
Foot contact is an important cue not only for human motion capture but also for motion understanding and physically plausible motion generation. However, most of the foot-contact annotations in existing datasets are estimated by purely visual matching and distance thresholding, which results in low accuracy and coarse granularity. Even though existing multimodal datasets synergistically capture plantar pressure (foot contact) and visual signals, they are specifically designed for small-range and slow motion such as Taiji Quan and Yoga. Therefore, there is still a lack of a vision-pressure multimodal dataset with large-range and fast human motion, as well as accurate and dense foot-contact annotation. To fill this gap, we propose a Multimodal MoCap Dataset with Vision and Pressure sensors, named MMVP. MMVP provides accurate and dense plantar pressure signals synchronized with RGBD observations, which is especially useful for both plausible shape estimation, robust pose fitting without foot drifting, and accurate global translation tracking. To validate the dataset, we propose an RGBD-P SMPL fitting method and also a monocular-video-based baseline framework, VP-MoCap, for human motion capture. Experiments demonstrate that our RGBD-P SMPL Fitting results significantly outperform pure visual motion capture. Moreover, VP-MoCap outperforms SOTA methods in foot-contact and global translation estimation accuracy. We believe the configuration of the dataset and the baseline frameworks will stimulate the research in this direction and also provide a good reference for MoCap applications in various domains. Project page: https://haolyuan.github.io/MMVP-Dataset/
@@ 1 Introduction
Human motion capture is an important foundation for motion analysis, behavior understanding, and pose generation, with a wide range of applications in AR/VR, disease diagnosis, robot manipulation, sports training etc. In recent years, human motion capture based on computer vision technology has been developing rapidly, which has mainly benefited from the development of deep learning technology as well as, more importantly, the release of a large number of human motion capture datasets. Most of the existing datasets focus on human shape and pose annotations. Given pairs of image and shape and pose annotations for training, learning-based methods can infer plausible human motion from visual signals alone [23, 5, 27, 25, 47, 59, 38, 26, 31, 12]. However, there are still many limitations when only using pose and shape annotations for supervision, resulting in artifacts such as foot drifting and erroneous global translation&rotation, etc.
@@
To solve the above problem, recent datasets introduce additional annotations, mainly contacts, such as foot-contact, body-scene contact [15, 18] or self-contact [10, 35], to enhance additional physical constraints for visual motion capture. Among the many additional contact annotations, the most classical one is the foot-contact annotation. With foot contact annotation, Rempe et al. [42] has inferred the 0-1 foot-contact label for the foot joints from an image, which successfully restricts the foot drifting artifacts and realizes much more accurate global translation results. It is worth mentioning that the labeled contact information can not only improve the accuracy of motion capture but also be used as important cues to directly improve the rationality of motion analysis [45, 46, 55] and even motion generation [16].
However, it is very difficult to accurately annotate the contact. Most of the existing datasets use purely visual matching for contact annotation  [15, 14, 18]. Specifically, these datasets usually require a pre-scanning of the 3D scene, and then, based on the image observations, use an optimization method to fit the parameterized human model (SMPL [33] or SMPL-X [40]) to the 3D scene, and finally the GT (ground truth) contact was annotated by thresholding the distance between the vertices of the SMPL model and the 3D scene. Since this is an in-direct annotation strategy, the accuracy is inevitably affected by the pose estimation error, shape estimation error, and even scene scanning error. The situation becomes even worse under fast and large-range motion when the contact time is short and the foot motion is sophisticated. In short, when the purely visual fitting error is larger, it is impractical to set a fixed distance threshold for accurate and detailed foot contact annotation. As a result, the foot contact annotation of these datasets is not precise enough and the annotation granularity is relatively coarse.
@@
As a result, there is still a lack of a motion capture dataset that can simultaneously take into account a large range of motion, fast body movements, and visual-pressure synergistic acquisition. To fill this gap, we design a vision-pressure synchronized multimodal human motion capture system and acquire a new dataset named Multimodal MoCap dataset with Vision and Pressure Sensors (MMVP), as shown in Fig. 1. Specifically, we captured single-view RGBD sequences and plantar pressure under a large range of rapid movements with sophisticated footwork. Based on the high-end insole sensor, we can obtain precise and dense foot pressure and contact annotations which are significantly more accurate and detailed than previous datasets. To fully utilize the synchronized RGBD and dense foot-contact signals for motion capture, we further design a novel RGBD-P SMPL fitting method based on the multimodal input, which is far more accurate than vision-only methods, especially around foot regions. Finally, based on the new dataset and the fitted SMPLs, we propose and evaluate a baseline framework for monocular video-based human motion capture, VP-MoCap, which incorporates a per-vertex-level foot-pressure predictor (FPP-Net) and contact-based pose optimization strategy together to significantly improve the global translation and pose estimation accuracy.
@@
  • We present MMVP, a novel human motion capture dataset that provides the most accurate foot pressure and dense contact annotations for large-range and rapid movements with detailed footwork.
  • An RGBD-P SMPL fitting method is introduced to fully utilize the multimodal signals and achieve high-quality human motion capture results with plausible shape, stable foot movement and accurate global translations. A monocular human MoCap baseline framework, VP-MoCap, consists of an FPP-Net and a joint optimization strategy for robust vision-based motion capture.
@@ 2.1 Multimodal MoCap Datasets
The Human datasets [19, 3, 1, 34, 41, 58] have greatly promoted the development of pose detection. Nevertheless, these datasets only include visual signals and provide annotations for body shapes and poses. To further analyze the human motion, multimodal datasets are needed. Tab. 1 summarizes existing multimodal datasets.
In terms of annotations, some datasets have started to incorporate annotations for the interactions, such as human-scene contact [15, 18], self-contact [10, 35] or object contact [6], providing new insights for the study and analysis of human motion. These contact information can significantly reduce pose ambiguities in motion capture and enhance the understanding of visual images. While these labels, whether manually annotated or based on distance thresholds, are still derived solely from visual information. Thus their accuracy is difficult to guarantee.
In terms of sensors, LiDAR is typically applied in large scenarios, and its signal modality is similar to vision [29, 8, 54, 54]. Inertial Measurement Units (IMUs) can measure the accelerations and orientations of body parts. Thus the combination of vision and IMUs is able to solve heavy occlusion or extreme illumination [51, 13]. In human gait analysis, clinical rehabilitation assessment and other biomedical fields, pressure is a commonly used signal. But medical datasets [11, 43, 52] often lack visual information. TMM100 [44] and MOYO [50] are the most representative vision & pressure datasets. The pressure can provide accurate contact information and rich physical details. However, both of these tend to favor slow and steady movements. Currently, there is still a lack of datasets that synchronize pressure and visual data to describe a wide range of fast body movements.
@@ 2.2 Human Pose Estimation
For distant targets, many current method utilize LIDAR to help global localization in mocap. LiDARCap [29] and HSC4D [7] employ LIDAR data to infer human 3D keypoints and obtain motion capture results. Similarly, SLOPER4D [8] utilizes LIDAR to perceive dynamic scene information for global human pose estimation. In order to tackle the lack of Off-grounded actions dataset, cimi4d [54] leverages LiDAR to acquire the human-scene contact information and apply a blending optimization process to regress human pose.
Several approaches have explored the fusion of RGB data with IMU data to enhance accuracy and robustness in the field of motion capture(mocap). Pan et al. [36] proposed a method that fuses RGB and IMU data using a complementary filter algorithm within the mocap framework, which addresses the challenge of reliability in vision information by incorporating information from the IMU. Matthew et al. [51] incorporates Long Short-Term Memory(LSTM) networks to fuse IMU tracking data and visual pose embeddings to improve the accuracy of pose estimation. HybridFusion [64] utilizes IMU data to address the occlusion problems in real-time monocular mocap system. EgoLocate [56] leverages mocap priors from inertial inputs to enhance the accuracy and robustness of the mocap system.
Thanks to the contact-based datasets, Contact-based methods [15, 14, 18, 10, 35] can estimate contact information from only images. Contact helps these method to reduce the ambiguity and achieve higher accuracy of pose estimation, which not only reduces ambiguity in pose estimation but also provides dynamic constraints for rigid body dynamics-related methods [45, 46, 55]. Furthermore, pressure can be utilized to embed physical constraints into pose estimation [60, 50].
@@ 3.1 Data Collection and Pre-processing
We collect a multimodal dataset containing synchronized RGBD video and foot pressure data. For visual data, we record the RGBD video using an Azure Kinect camera111[Online].https://azure.microsoft.com/en-us/products/kinect-dk/at a frame rate of 30Hz. For pressure data, we use Xsensor pressure insoles (HX 210-510)222[Online].https://www.xsensor.com/solutions-and-platform/human-performance/gait-motion-insoles. Each Xsensor pressure insole contains 242 independent pressure sensors that accurately capture the changes in pressure, with a frame rate of up to 150Hz. These two data are collected together. Due to the hardware limitation, we cannot synchronize the two streams of signals automatically, so we do manual synchronization before usage. The MMVP dataset covers up to 10 types of large-range and fast movements, including running, standing-long-jumping, and skipping, among other common types of exercises and dances. The majority of the subjects are teenagers, covering different age groups and both genders.
@@ 3.2 Calculate Dense Foot Contact
Benefiting from the highly accurate pressure insole, we can calculate precise and dense foot contact annotations. To ease the usage of the dense pressure and contact information, we define it on the surface of the popular SMPL model [33]. We select 192 vertices in total on the feet of the SMPL model and the dense contact label can be defined as $C\in\mathbb{R}^{192}$.
A typical method to annotate the dense contact is to directly thresholding the pressure. However, the foot pressure can vary dramatically depending on body weight and motion status, and a fixed value threshold is hard to design. Luckily, with insoles, it’s easy for us to get the accurate body weight of different subjects by simply summarizing the pressure at the beginning when the subject is standing still. Therefore we normalized and mapped the pressure to $[0,1]$ as
@@
Figure 4: VP-MoCap pipeline. Given a RGB sequence, RTMPose [20] and CLIFF [31] are applied to detect 2D key points and regress initial pose. FPP-Net predicts foot pressure distribution and dense foot contact with keypoint sequence. Guided by foot contact, joint optimization is applied to estimate pose and trajectory. (Green represents the learning part, while orange represents the optimization part.)
@@ 3.3 RGBD-P SMPL Fitting
As illustrated in Fig. 2, existing datasets calculate contact after fitting body model to pre-scaned scene with purely vision signals. Therefore, it is difficult to accurately maintain the interaction relationship between human body and scene during fitting. By constructing loss terms according to the ready-to-use contact information derived from the pressure signal, we can obtain more accurate pose and shape fitting results. Specifically, we use SMPL [33] to numerically represent the body shape and pose. The SMPL model takes pose parameters $\boldsymbol{\theta}\in\mathbb{R}^{72}$, shape parameters $\boldsymbol{\beta}\in\mathbb{R}^{10}$, global rotation $\boldsymbol{R}\in SO(3)$, and global translation $\boldsymbol{T}\in\mathbb{R}^{3}$ as input, producing a triangulated mesh of 6890 vertices. The $k$ joints of the model are represented as $\boldsymbol{J}(\boldsymbol{\theta},\boldsymbol{\beta},\boldsymbol{R},\boldsymbol{T})\in\mathbb{R}^{k\times 3}$. To guarantee the shape-fitting performance for teenagers with various body scales in our dataset, we follow the operation in AGORA dataset [37], which introduces a blending parameter $\alpha$ to balance the shape difference between children and adults. According to AGORA, SMPL template $\boldsymbol{M}_{A}$ and SMIL [17] template $\boldsymbol{M}_{C}$ are blended as $M_{blend}=\alpha\boldsymbol{M}_{A}+(1-\alpha)\boldsymbol{M}_{C}$. The vertices of the blended model are denoted as $\boldsymbol{V}$.
@@
where $\Pi_{floor}({\boldsymbol{V}_{foot}})$ denotes the distance between ground and human foot surface vertex whose contact label is 1.
@@ 4 Method: VP-MoCap
Previous monocular RGB based human mesh recovery methods [31] usually suffers from translation drifting, foot sliding, and foot-plane penetration, which make them unstable when handling relatively large-range and rapid movements. In the RGBD-P SMPL fitting section, we observed a much more accurate and plausible pose and translation by using depth and pressure signals for joint optimization. Thus, we further explore the question that: given plausible dense contact and 3D scene information inferred from the RGB images directly, if we can also achieve stable and physically plausible human motion capture results for large-range and rapid movements.
Therefore, we first construct FPP-Net to estimate foot contact and then use the off-the-shelf monocular RGB depth estimation method to estimate depth map of the ground from a single view RGB video $\mathcal{V}=\{I_{1},I_{2},...I_{t}\}$. Finally, when provided with foot contact and ground depth, we try to obtain better pose and translation through optimization only through monocular RGB image sequence. The overall framework is summarized in Fig. 4.
@@ 4.1 FPP-Net
To estimate foot pressure distribution and foot contact from a RGB sequence, we design a foot pressure predictor (FPP-Net), which is illustrated in green dashed box in Fig. 4. Rather than using raw images to regress the final results, we use the 2D keypoints sequence $\mathcal{J}^{2d}=\{j^{2d}_{1},j^{2d}_{2},...j^{2d}_{t}\}$ detected by RTMPose [20] instead as input, as we assume that foot pressure and contact are more related to body pose and motion dynamic state. The task can be formulated as

### v2

@@ Abstract
Foot contact is an important cue for human motion capture, understanding, and generation. Existing datasets tend to annotate dense foot contact using visual matching with thresholding or incorporating pressure signals. However, these approaches either suffer from low accuracy or are only designed for small-range and slow motion. There is still a lack of a vision-pressure multimodal dataset with large-range and fast human motion, as well as accurate and dense foot-contact annotation. To fill this gap, we propose a Multimodal MoCap Dataset with Vision and Pressure sensors, named MMVP. MMVP provides accurate and dense plantar pressure signals synchronized with RGBD observations, which is especially useful for both plausible shape estimation, robust pose fitting without foot drifting, and accurate global translation tracking. To validate the dataset, we propose an RGBD-P SMPL fitting method and also a monocular-video-based baseline framework, VP-MoCap, for human motion capture. Experiments demonstrate that our RGBD-P SMPL Fitting results significantly outperform pure visual motion capture. Moreover, VP-MoCap outperforms SOTA methods in foot-contact and global translation estimation accuracy. We believe the configuration of the dataset and the baseline frameworks will stimulate the research in this direction and also provide a good reference for MoCap applications in various domains. Project page: https://metaverse-ai-lab-thu.github.io/MMVP-Dataset/
11footnotetext: Equal contribution.22footnotetext: Corresponding author.
@@ 1 Introduction
Human motion capture is an important foundation for motion analysis, behavior understanding, and pose generation, with a wide range of applications in AR/VR, disease diagnosis, robot manipulation, sports training, etc. In recent years, human motion capture based on computer vision technology has been developing rapidly, which has mainly benefited from the development of deep learning technology as well as, more importantly, the release of a large number of human motion capture datasets. Most of the existing datasets focus on human shape and pose annotations. Given pairs of image and shape and pose annotations for training, learning-based methods can infer plausible human motion from visual signals alone [23, 5, 27, 25, 47, 59, 38, 26, 31, 12]. However, there are still many limitations when only using pose and shape annotations for supervision, resulting in artifacts such as foot drifting and erroneous global translation&rotation, etc.
@@
To solve the above problem, recent datasets introduce additional annotations, mainly contacts, such as foot-contact, body-scene contact [15, 18] or self-contact [10, 35], to enhance additional physical constraints for visual motion capture. Among these contact annotations, the most classical one is the foot-contact annotation. With foot contact annotation, Rempe et al. [42] has inferred the 0-1 foot-contact label for the foot joints from an image, which successfully restricts the foot drifting artifacts and realizes much more accurate global translation results. It is worth mentioning that the labeled contact information can not only improve the accuracy of motion capture but also be used as important cues to directly improve the rationality of motion analysis [45, 46, 55] and even motion generation [16].
However, it is very difficult to accurately annotate the contact. Most of the existing datasets use purely visual matching and distance thresholding for contact annotation  [15, 14, 18]. Specifically, these datasets usually require a pre-scanning of the 3D scene, and then, based on the image observations, use an optimization method to fit the parameterized human model (SMPL [33] or SMPL-X [40]) to the 3D scene, and finally the GT (ground truth) contact was annotated by thresholding the distance between the vertices of the SMPL model and the 3D scene. Since this is an in-direct annotation strategy, the accuracy is inevitably affected by the pose estimation error, shape estimation error, and even scene scanning error. The situation becomes even worse under fast and large-range motion when the contact time is short and the foot motion is sophisticated. In short, when the purely visual fitting error is larger, it is impractical to set a fixed distance threshold for accurate and detailed foot contact annotation. As a result, the foot contact annotation of these datasets is not precise enough and the annotation granularity is relatively coarse.
@@
As a result, there is still a lack of a motion capture dataset that can simultaneously take into account a large range of motion, fast body movements, and visual-pressure synergistic acquisition, as well as accurate and dense foot-contact annotation. To fill this gap, we design a vision-pressure synchronized multimodal human motion capture system and acquire a new dataset named Multimodal MoCap dataset with Vision and Pressure Sensors (MMVP), as shown in Fig. 1. Specifically, we captured single-view RGBD sequences and plantar pressure under a large range of rapid movements with sophisticated footwork. Based on the high-end insole sensor, we can obtain precise and dense foot pressure and contact annotations which are significantly more accurate and detailed than previous datasets. To fully utilize the synchronized RGBD and dense foot-contact signals for motion capture, we further design a novel RGBD-P SMPL fitting method based on multimodal input, which is far more accurate than vision-only methods, especially around foot regions. Finally, based on the new dataset and the fitted SMPLs, we propose and evaluate a baseline framework for monocular video-based human motion capture, VP-MoCap, which incorporates a per-vertex-level foot-pressure predictor (FPP-Net) and contact-based pose optimization strategy together to significantly improve the global translation and pose estimation accuracy.
@@
  • We present MMVP, a novel multimodal human motion capture dataset that provides precise foot pressure and the most accurate dense contact annotations for large-range and rapid movements with detailed footwork.
  • An RGBD-P SMPL fitting method is introduced to fully utilize the multimodal signals and achieve high-quality human motion capture results with plausible shape, stable foot movement, and accurate global translations. We also introduce a monocular RGB-based human motion capture baseline framework called VP-MoCap, which combines an FPP-Net and a joint optimization strategy.
@@ 2.1 Multimodal MoCap Datasets
The Human datasets [19, 3, 1, 34, 41, 58] have greatly promoted the development of pose detection. Nevertheless, these datasets only include visual signals and provide annotations for body shapes and poses. To further analyze human motion, multimodal datasets are needed. Tab. 1 summarizes existing multimodal datasets.
Some datasets introduce additional interaction annotations, such as human-scene contact [15, 18], self-contact [10, 35] or object contact [6], providing new insights for the study and analysis of human motion. These contacts can significantly reduce pose ambiguities in motion capture and enhance the understanding of visual images. Whether manually annotated or based on thresholds, these labels are derived solely from visual information, making it difficult to guarantee accuracy.
Some datasets incorporate signals from other modalities. LiDAR is typically applied in large scenarios, and its signal modality is similar to vision [29, 8, 54, 54]. Inertial Measurement Units (IMUs) can measure the accelerations and orientations of body parts. Thus the combination of vision and IMUs can solve heavy occlusion or extreme illumination [51, 13]. Pressure is a commonly used signal in human gait analysis, clinical rehabilitation assessment, and other biomedical fields. But medical datasets [11, 43, 52] often lack visual information. TMM100 [44] and MOYO [50] are the most representative vision & pressure datasets. The pressure can provide accurate contact information and physical features. However, both of these tend to favor slow and steady movements. Currently, there is still a lack of datasets that synchronize pressure and visual data to describe a wide range of fast body movements.
@@ 2.2 Human Pose Estimation
LIDAR is employed to localize the global position of distant targets. LiDARCap [29], HSC4D [7], and Sloper4d [8] employ LIDAR data to estimate global human poses. CIMI4D [54] leverages LiDAR to acquire the human-scene contact information and regress the human pose for off-grounded motions. Due to the distance of the subjects, these methods often lack fine-grained motion details.
Some approaches have explored the fusion of RGB data with IMU data to enhance the accuracy and robustness of pose estimation. Matthew et al. [51] incorporates Long Short-Term Memory (LSTM) networks to fuse IMU tracking data and visual pose embeddings to improve the accuracy of pose estimation. HybridFusion [64] utilizes IMU data to address the occlusion problems in a real-time monocular MoCap system. Pan et al. [36] fuses RGB and IMU data using a complementary filter algorithm within the mocap framework to address the challenge of reliability in vision information. EgoLocate [56] leverages MoCap priors from inertial inputs to enhance the accuracy and robustness of the mocap system.
Contact provides strong interaction priors for pose estimation. Thanks to the contact-based datasets, many methods [15, 18, 10] can estimate contact from a single image. The contact helps not only reduce ambiguity and achieve higher accuracy in pose estimation [35, 14, 10] but also provides dynamic constraints for rigid body dynamics-related methods [45, 46, 55]. In addition to determining contact, pressure can also provide force information. Zhang et al. [60] estimate human pose from three depth cameras and a pair of pressure-sensing shoes. Tripathi et al. [50] enhance the stability in pose estimation by encouraging plausible floor contact and overlapping the center of pressure and the SMPL’s center of mass.
@@ 3.1 Data Collection and Pre-processing
We collect a multimodal dataset containing synchronized RGBD video and foot pressure data. For visual data, we record the RGBD video using an Azure Kinect camera at a frame rate of 30Hz. For pressure data, we use Xsensor pressure insoles (HX 210-510). Each Xsensor pressure insole contains 242 independent pressure sensors that accurately capture the changes in pressure, with a frame rate of up to 150Hz. Due to the hardware limitation, we cannot synchronize the two streams of signals automatically, so we do manual synchronization before usage. The MMVP dataset covers up to 10 types of large-range and fast movements, including running, standing-long-jumping, and skipping, among other common types of exercises and dances. The majority of the subjects are teenagers, covering different age groups and both genders. For more details, please refer to the supplementary material.
@@ 3.2 Calculate Dense Foot Contact
Benefiting from the highly accurate pressure insole, we can calculate precise and dense foot contact annotations. To ease the usage of the dense pressure and contact information, we define it on the surface of the SMPL [33]. We select 192 vertices in total on the feet of the SMPL model and the dense contact label can be defined as $C\in\mathbb{R}^{192}$.
The foot pressure can vary dramatically depending on body weight and motion status. So it’s hard to design a fixed-value threshold to annotate the dense contact. Luckily, with insoles, it’s easy for us to get the accurate body weight of different subjects by simply summarizing the pressure at the beginning when the subject is standing still. Therefore we normalized and mapped the pressure to $[0,1]$ as
@@
Figure 4: VP-MoCap pipeline. Given an RGB sequence, RTMPose [20] and CLIFF [31] are applied to detect 2D keypoints and regress the initial pose. FPP-Net predicts foot pressure distribution and dense foot contact with keypoint sequence. Guided by foot contact, joint optimization is applied to estimate pose and trajectory. (Green represents the learning part, while orange represents the optimization part.)
@@ 3.3 RGBD-P SMPL Fitting
As illustrated in Fig. 2, existing datasets calculate contact after fitting the body model to the pre-scanned scene with purely vision signals. Therefore, it is difficult to accurately maintain the interaction relationship between the human body and the scene during fitting. By constructing loss terms according to the ready-to-use contact information derived from the pressure signal, we can obtain more accurate pose and shape-fitting results. Specifically, we use SMPL [33] to numerically represent the body shape and pose. The SMPL model takes pose parameters $\boldsymbol{\theta}\in\mathbb{R}^{72}$, shape parameters $\boldsymbol{\beta}\in\mathbb{R}^{10}$, global rotation $\boldsymbol{R}\in SO(3)$, and global translation $\boldsymbol{T}\in\mathbb{R}^{3}$ as input, producing a triangulated mesh of 6890 vertices. The $k$ joints of the model are represented as $\boldsymbol{J}(\boldsymbol{\theta},\boldsymbol{\beta},\boldsymbol{R},\boldsymbol{T})\in\mathbb{R}^{k\times 3}$. To guarantee the shape-fitting performance for teenagers with various body scales in our dataset, we follow the operation in AGORA dataset [37], which introduces a blending parameter $\alpha$ to balance the shape difference between children and adults. According to AGORA, SMPL template $\boldsymbol{M}_{A}$ and SMIL [17] template $\boldsymbol{M}_{C}$ are blended as $M_{blend}=\alpha\boldsymbol{M}_{A}+(1-\alpha)\boldsymbol{M}_{C}$. The vertices of the blended model are denoted as $\boldsymbol{V}$.
@@
where $\Pi_{floor}({\boldsymbol{V}_{foot}})$ denotes the distance between the ground and the foot surface vertex in contact.
@@ 4 Method: VP-MoCap
Previous monocular RGB-based human mesh recovery methods [31] usually suffer from translation drifting, foot sliding, and foot-plane penetration, which make them unstable when handling relatively large-range and rapid movements. In the RGBD-P SMPL fitting section, we observed a much more accurate and plausible pose and translation by using depth and pressure signals for joint optimization. Thus, we further explore a more general question: given only monocular RGB sequences, can we also achieve stable and physically plausible human motion capture results for large-range and rapid movements with inferred dense contact and 3D scene?
Therefore, we first construct FPP-Net to estimate foot contact and then use Zoedepth [2] to estimate the 3D ground from a monocular RGB sequence. Finally, with foot contact and ground depth, we try to obtain better pose and translation through optimization using only the monocular RGB sequence. The overall framework is summarized in Fig. 4.
@@ 4.1 FPP-Net
To estimate foot pressure distribution and foot contact from an RGB sequence, we design a foot pressure predictor (FPP-Net), which is illustrated in the green dashed box in Fig. 4. We assume that foot pressure and contact are more related to body pose and motion dynamic state. Thus we follow Physcap [45] and Jesse et al. [44] to use sequential 2D keypoints for contact prediction. The goal of this strategy is to also decouple the network training process from the dataset-capturing environments (when compared with direct image regression methods like BSTRO [18] and DECO [49]) thus significantly enhancing the generalization capacity. The 2D keypoints sequence $\mathcal{J}^{2d}=\{j^{2d}_{1},j^{2d}_{2},...j^{2d}_{t}\}$ is detected by RTMPose [20]. The task can be formulated as